OpenStudy (anonymous):

ok guys last problem and I won't bother you anymore! a business purchases a van for $27,000. After 5 years it's depreciated value will be $12,000. write an equation, find the value after 2 years

OpenStudy (amistre64):

27000 - some sort of linear depreciation for 5 years = 12000

OpenStudy (amistre64):

soo; 12000 = 27000 - 5(dep per year)

OpenStudy (amistre64):

dep per year is then: (12000-27000)/-5

OpenStudy (amistre64):

does that make sense?

OpenStudy (anonymous):

lol kinda so in the book t= time and v = the value of the van in terms of years

OpenStudy (amistre64):

v(t) = original price - dep(time) v(t) = 27000 - dep(t) where dep is determined fromthe givein information

OpenStudy (amistre64):

27000 -12000 -------- 15000 over 5 years; is just 3000 per year

OpenStudy (amistre64):

v(t) = 27000 - 3000(t) when t=2 we get ?

OpenStudy (anonymous):

6000 if its 3000 a year right?

OpenStudy (amistre64):

correct, sooo 27000 -6000 ------ van value

OpenStudy (anonymous):

the example in the book gives like 100 extra steps

OpenStudy (amistre64):

my arthritis doesnt let me do 100 steps :)

OpenStudy (amistre64):

we could try to poit this to get a slope ... (0,27000) and (5,12000) 0,27000 -5,-12000 ---------- -5, 15000; slope = 15000/-5 = -3000 y-(27000) = -3000(x-0) y = -3000x + 27000

OpenStudy (amistre64):

ouch ....

OpenStudy (amistre64):

i find steps to be pointless when compared to reason ...

OpenStudy (anonymous):

yay now my homework is finished, i can scan it, send it in, and get back to fixing my breaks:( i was supposed to get to school today but I can't without breaks:( and if you don't mind me asking, how old are you?